OpenStudy (anonymous):

Find the sum of the expression 6/11+ (−3/8) ) Circle your final answer.

OpenStudy (anonymous):

Try finding the common denominator first.

OpenStudy (anonymous):

6

OpenStudy (anonymous):

So, first thing's first. You should find a common denominator, which is a number that both 11 and 8, your denominators, are divisible by. The easiest way to do this would be to multiply 11 and 8, giving you what?

OpenStudy (anonymous):

88

OpenStudy (anonymous):

Yep! So in order to make it nice and easy to add, then multiply the numerator and denominator by the same number.\[(\frac{ 6 }{ 11 } * \frac{ 8 }{ 8 }) - (\frac{ 3 }{ 8 } * \frac{ 11 }{ 11 }) = ?\]

OpenStudy (anonymous):

You're able to do this because 8/8 and 11/11 equals 1, so you're not changing the value of the numbers! You're simply doing the opposite of reducing them, in order to make it easier to subtract.

OpenStudy (anonymous):

ok ok